OpenStudy (anonymous):

What is 15 degrees Celsius (°C) in degrees Fahrenheit (°F)? Use the formula F = 1.8C + 32.

OpenStudy (anonymous):

F = 1.8*15 + 32 F = 27 + 32 F = 59 degrees Fahrenheit.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Well done, you solved your own question :)